OpenStudy (toxicsugar22):

give the exact coordinates of any four points on the hyperbola (y+2)^2/6-(x+2)^2/10=1

OpenStudy (anonymous):

Just plug in any value for x, then solve for y...

OpenStudy (toxicsugar22):

can u help me

OpenStudy (toxicsugar22):

please so i know how to do it

OpenStudy (anonymous):

okay, you have \[((y+2)^{2})/6 - ((x-2)^{2})/10 = 1\] \[((y+2)^{2})/6 - ((3-2)^{2})/10 = 1\] now solve for y \[((y+2)^{2})/6 - 1/10 = 1\]

OpenStudy (toxicsugar22):

give the exact coordinates of any four points on the hyperbola (y+2)^2/6-(x+2)^2/10=1

OpenStudy (anonymous):

I showed you how to solve this above.

OpenStudy (toxicsugar22):

yes so now solve for y

OpenStudy (toxicsugar22):

can u help me

OpenStudy (toxicsugar22):

can u